a ball rolls horizontally off the top of a stairway with a speed of 1.52m/s . the steeps are 20.3m high and 20.3 m wide which step does the ball hit first?

Grade:11

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear shivanshi

its a question of projectile motion

initial velocity of ball is 1.52 m/sec in horizontal direction

the steeps are 20.3m high

 time take to fall this distance is

  h= 1/2gt2

 20.3 =1/2 *10 *t2

 t= 2.015 sec

in this time horizontal distane travell = 1.52 * 2.015 =3.063 m

since distance is less than 20.3  so ball hit the first step

Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin